# InstructTA: Instruction-Tuned Targeted Attack for Large Vision-Language Models

### v2

@@ Abstract
Large vision-language models (LVLMs) have demonstrated their incredible capability in image understanding and response generation. However, this rich visual interaction also makes LVLMs vulnerable to adversarial examples. In this paper, we formulate a novel and practical gray-box attack scenario that the adversary can only access the visual encoder of the victim LVLM, without the knowledge of its prompts (which are often proprietary for service providers and not publicly available) and its underlying large language model (LLM). This practical setting poses challenges to the cross-prompt and cross-model transferability of targeted adversarial attack, which aims to confuse the LVLM to output a response that is semantically similar to the attacker’s chosen target text. To this end, we propose an instruction-tuned targeted attack (dubbed InstructTA) to deliver the targeted adversarial attack on LVLMs with high transferability. Initially, we utilize a public text-to-image generative model to “reverse” the target response into a target image, and employ GPT-4 to infer a reasonable instruction $\boldsymbol{p}^{\prime}$ from the target response. We then form a local surrogate model (sharing the same visual encoder with the victim LVLM) to extract instruction-aware features of an adversarial image example and the target image, and minimize the distance between these two features to optimize the adversarial example. To further improve the transferability, we augment the instruction $\boldsymbol{p}^{\prime}$ with instructions paraphrased from an LLM. Extensive experiments demonstrate the superiority of our proposed method in targeted attack performance and transferability.
@@ 1 Introduction
Large vision-language models (LVLMs) have achieved great success in text-to-image generation [35], image captioning [7] and visual question-answering [25], due to the colossal increase of training data and model parameters. Benefiting from the strong comprehension of large language models (LLMs), recent LVLMs [25, 12, 48, 27] on top of LLMs have demonstrated a superior capability in solving complex tasks than that of small vision-language models. It is noted that GPT-4 [32] supports visual inputs and perform human-like conversations in various complicated scenarios.
@@
However, recent studies [46, 15, 36, 3] exposed that LLMs with visual structures could be easily fooled by adversarial examples (AEs) crafted via adding imperceptible perturbations to benign images. In particular, targeted attack with malicious goal raises more serious safety concerns to LVLMs and downstream applications. For example, targeted adversarial examples could jailbreak LVLMs to explain the bomb-making process or trick medical diagnostic models into generating reports of specific diseases for insurance. It is thus demanding to comprehensively study adversarial attacks on LVLMs to address these security threats before security breaches happen.
@@
This paper focuses on the targeted attack on LVLMs, where the adversary aims to generate an adversarial image that misleads the LVLM to respond with a specific target text, despite varying instructions the LVLM may use. We target a practical gray-box setting, where the adversary can only access the visual encoder of the LVLM without any knowledge of the LLM and instructions leveraged by the LVLM. [^1] To clarify, in this paper, we use the term “instruction” and “prompt” interchangeably. While de facto LVLMs often leverage standard visual encoders that are publicly known, the employed instructions are often proprietary for service providers and not publicly available. This practical assumption brings some key challenges: (1) The attacker cannot easily gain the instruction-aware features from the accessed visual encoder to launch exact targeted attacks due to a lack of instructions. (2) Cross-model and cross-instruction transferability places new demands on the targeted attack, since the target LVLM may have distinct LLM backends and use varying (unknown) instructions.
@@
Figure 1: The framework of our instruction-tuned targeted attack (InstructTA). Given a target text $\boldsymbol{y}_{t}$, we first transform it into the target image $\boldsymbol{x}_{t}$ with a text-to-image model $h_{\xi}$. Simultaneously, GPT-4 infers a reasonable instruction $\boldsymbol{p}^{\prime}$. When feeding the augmented instruction $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ rephrased by an LLM $\boldsymbol{G}_{3.5}$, the surrogate model $M$ extracts instruction-aware features of the AE $\boldsymbol{x}^{\prime}$ and $\boldsymbol{x}_{t}$, respectively. Finally, we minimize the $L_{2}$ distance between these two features to optimize the AE $\boldsymbol{x}^{\prime}$. Notably, the instruction $\boldsymbol{p}_{i/j}^{\prime}$ is updated on-the-fly during the optimization of the adversarial sample.
@@
  • This paper explores the attack vector of LVLMs under a practical yet unexplored gray-box setting, i.e., the adversary only has the rights to access the visual encoder in LVLMs, excluding the LLM and instructions.
  • We propose an effective targeted attack, InstructTA. It cleverly leverages text-to-image models and LLMs to infer reasonable target images and instructions, respectively, and augments the transferability of adversarial samples with instruction-tuning paradigms.
@@ 2 Related Work
LVLMs. Benefiting from the powerful capabilities of LLMs in answering diverse and complex questions [31], LVLMs [46] have demonstrated remarkable performance on a variety of multi-modal tasks, e.g., image captioning, visual question-answering, natural language visual reasoning and visual dialog. To bridge the image-text modality gap, BLIP-2 [25] proposes the Q-Former to extract visual features into the frozen OPT [45] or FlanT5 [11]. MiniGPT-4 [48] employs the identical pretrained ViT [16] and Q-Former as BLIP-2, but uses the more powerful Vicuna [10] to serve as the base LLM for text generation. InstructBLIP [12] introduces an instruction-aware Q-Former to extract the most task-relevant visual representations. LLaVA [27] bridges a pre-trained visual encoder with LLaMA, and fine-tunes them with question-image-answer paired data generated by GPT-4.
@@
Adversarial Attacks. Although adversarial attacks are now being developed in a variety of modalities and tasks, the earliest and most extensive research occurred in the field of image classification [4, 37]. In this case, the goal of adversarial attack is often to fool a well-trained visual classifier by crafted adversarial images. Based on their purpose, we divide adversarial attacks into targeted attacks and non-targeted attacks [2]. Targeted attacks aim to confuse the model to recognize the adversarial example as a target label, whereas non-targeted attacks only ensure that the model make wrong predictions, not caring about the output error category. Moreover, according to the information of the attacked model exposed to the adversary, adversarial attacks can be grouped into white-box attack and black-box attack. The white-box attacks can obtain the whole parameters and structure of the target model. Thus, they generally optimize adversarial perturbations with gradients of loss w.r.t inputs, e.g., FGSM [19], I-FGSM [22], C&W[5] and PGD [30]. In contrast, only the input and output of the attacked model can be captured by black-box attack, which is more challenging and practical than white-box one. One solution is to estimate the gradient or search for successful perturbations by directly querying the target model multiple times, e.g., ZOO [8] and NES attack [21]. Another way is to implement transfer attacks by constructing adversarial samples from substitute models, e.g., SBA [33] and ensemble-based approaches [28, 14, 18, 20].
@@
AEs on LVLMs. Currently, there are some pioneering and contemporaneous works of targeted attacks on LVLMs. On one hand, some works aim to mislead the LVLMs to output specific words. For example, Bailey et al. [3] forces the LVLM to respond with the target string in a white-box setting. Dong et al. [15] present text description attack to mislead the black-box Bard [1] to output targeted responses, which maximize the log-likelihood of the substitute LVLMs to predict the target sentence. On the other hand, some works have been devoted to fooling LVLMs into outputting text with the same semantics as the target. For example, Zhao et al. [46] propose three feature matching methods (i.e., MF-it, MF-ii and MF-tt) to launch targeted attack with given target text against LVLMs in black-box settings. Shayegani et al. [36] also align the adversarial image and target images in the embedding space for LVLM jailbreak. Different from [46], we focus on the gray-box setting instead of the black-box setting, and further improve the attack transferability under the guidance of inferred and rephrased instructions. While [36] also accesses visual encoders in LVLMs for jailbreak, our work highlights the capability to implement generic targeted attacks even without knowing the target LVLM’s instructions.
@@ 3.1 Problem Formulation
LVLM. Given an image $\boldsymbol{x}$ and a text instruction $\boldsymbol{p}$ as inputs, the LVLM $F$ outputs an answer $\boldsymbol{y}$. Formally,
@@
where $F_{\theta}$ denotes the LVLM parameterized by $\theta$. Typically, the LVLM [25, 48, 12, 27] projects the extracted visual features into the word embedding space and then feeds them to its underlying LLM. Hence, LVLM consists of a pre-trained visual encoder, a projector and a well-trained LLM, as shown in 2.
@@
where $\boldsymbol{x}$ is a given benign image sample, $\boldsymbol{x}^{\prime}$ represents the adversarial example of $\boldsymbol{x}$, and $\boldsymbol{y}_{t}$ denotes the predefined target text. $\epsilon$ is the budget of controlling adversarial perturbations, and $\boldsymbol{p}$ denotes the instructions used by the target LVLM ($\boldsymbol{p}$ is unknown to attackers; see § 3.2). $S(\cdot,\cdot)$ is a metric depicting the semantic similarity between two input sentences. This way, Eqn. 2 ensures that the response of the adversarial example shall retain the same meaning as the target text $y_{t}$.
@@
0:  a benign image $\boldsymbol{x}$, a target text $\boldsymbol{y}_{t}$, a text-to-image generative model $h_{\xi}$, a surrogate model $M$, a VLP model with the visual encoder $f$ and text encoder $g$, ChatGPT $G_{3.5}$, GPT-4 $G_{4}$, custom prompts $\boldsymbol{p}_{g}$ and $\boldsymbol{p}_{r}$, step size $\eta$, perturbation budget $\epsilon$, attack iterations $T$
1:  Initialize: $\boldsymbol{x}^{\prime}=\boldsymbol{x}$, $\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$, $\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t})$  // Eqn. (3)
@@
3:     $\boldsymbol{p}_{i}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
4:     $\boldsymbol{p}_{j}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
@@
Figure 3: An example of rephrasing an instruction. Given the target response $\boldsymbol{y}_{t}$, GPT-4 infers an instruction, i.e., “Can you describe the scene depicted in the photograph?”. The real instruction assigned to this target text is “Explain the visual content of the image in great detail.”
3.2 Targeted Attack with Instruction Tuning
Assumption. In this paper, we consider a practical gray-box setting, where the adversary only knows the visual encoder used in the attacked model, excluding LLMs and instructions of the victim LVLM. That is, the adversary is agnostic to $\boldsymbol{p}$. Instead, we assume that attackers know the usage domain of the target LVLM, e.g., CT image diagnosis; considering $F$ is often hosted as a task-specific service, attackers can easily obtain this information. Accordingly, attackers can locally prepare instructions $\boldsymbol{p}^{\prime}$ which are commonly used in the target domain (see below for details). We also assume that before attack, attackers can form a local surrogate model with the same visual encoder as that of the victim model.
Inferring Instructions. We first locally infer the possible instructions $\boldsymbol{p}^{\prime}$ for the target text $\boldsymbol{y}_{t}$ by using GPT-4 [32]. Formally,
$\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t}),$ (3)
where $G_{4}$ denotes the language-only GPT-4, and $\boldsymbol{p}^{\prime}$ represents an instruction generated by $G_{4}$. The above action denotes that we feed an instruction $\boldsymbol{p}_{g}$ concatenated with the target response text $\boldsymbol{y}_{t}$ to the GPT-4. In the experiment, we set $\boldsymbol{p}_{g}$ to a straightforward prompt “What is the question or prompt for this answer: ”.
Forming Surrogate Models. Once the instructions have been inferred, the surrogate model uses a projector to transform the features from the visual encoder into instruction-aware features. Thus, the surrogate model comprises two modules: the visual encoder from the victim model and the pre-trained projector (e.g., Q-Former [25]). The projector maps global visual features into informative features tailored to given instructions.
Augmenting Transferability. Our tentative exploration shows that AEs generated by Eqn. 4, although manifesting reasonable attackability, are often less transferable to the wide range of possible instructions in the wild. Thus, to improve the transferability and robustness of the generated adversarial examples, we use ChatGPT (i.e., GPT-3.5) to rephrase the inferred $\boldsymbol{p}^{\prime}$ as augmentation. The paraphrase prompt template fed into the ChatGPT is “Paraphrase this sentence: ”. This way, Eqn. 4 can be rewritten as
$\min_{\boldsymbol{x}^{\prime}}\mathcal{J}_{ins}=\mathbb{E}_{\boldsymbol{p}_{i}^{\prime},\boldsymbol{p}_{j}^{\prime}}\|M(\boldsymbol{x}^{\prime}$ $,\boldsymbol{p}_{i}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}_{j}^{\prime})\|_{2}^{2},$ (5)
[…]
where $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ are augmented instructions from $\boldsymbol{p}^{\prime}$. Both of them are generated by ChatGPT, i.e., $\boldsymbol{p}_{i}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, $\boldsymbol{p}_{j}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, where $G_{3.5}$ represents ChatGPT and $\boldsymbol{p}_{r}$ denotes the aforementioned rephrasing template. As illustrated in Fig. 3, ChatGPT is generally effective to rephrase a seed instruction $\boldsymbol{p}^{\prime}$ into a set of semantically-correlated instructions.
Table 1: Targeted attacks against victim LVLMs. A total of 1,000 clean images denoted as $\boldsymbol{x}$ are sampled from the ImageNet-1K [13] validation set. For each of these benign images, a targeted text $\boldsymbol{y}_{t}$ and a prompt $\boldsymbol{p}$ are randomly assigned from LLaVA-Instruct-150K [27]. Our analysis involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on attacked LVLMs, which was performed using several CLIP text encoders and their ensemble. We also report the No. of successful attacks within 1,000 adversarial samples, determined by GPT-4. CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.
VLVM Attack method Text encoder (pretrained) for evaluation The No. of successes RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.296 0.427 0.342 0.354 0.171 0.318 4 MF-it (CLIP) [46] 0.292 0.424 0.339 0.353 0.168 0.315 2 MF-ii (CLIP) [46] 0.293 0.426 0.340 0.353 0.168 0.316 2 MF-it (EVA-CLIP) 0.574 0.667 0.616 0.630 0.481 0.594 519 MF-ii (EVA-CLIP) 0.573 0.661 0.616 0.627 0.484 0.592 514 InstructTA-MF-it (ours) 0.604 0.691 0.644 0.656 0.515 0.622 652 InstructTA-MF-ii (ours) 0.599 0.682 0.639 0.649 0.512 0.616 619 InstructBLIP Clean image 0.392 0.557 0.420 0.421 0.312 0.420 0 MF-it (CLIP) [46] 0.394 0.558 0.423 0.422 0.313 0.422 0 MF-ii (CLIP) [46] 0.395 0.557 0.424 0.423 0.314 0.422 0 MF-it (EVA-CLIP) 0.773 0.822 0.791 0.796 0.740 0.785 292 MF-ii (EVA-CLIP) 0.761 0.813 0.780 0.783 0.727 0.773 259 InstructTA-MF-it (ours) 0.810 0.849 0.826 0.830 0.784 0.820 383 InstructTA-MF-ii (ours) 0.789 0.833 0.805 0.809 0.758 0.799 328 MiniGPT-4 Clean image 0.345 0.495 0.367 0.374 0.253 0.367 0 MF-it [46] 0.341 0.493 0.364 0.369 0.251 0.364 0 MF-ii [46] 0.340 0.492 0.363 0.369 0.249 0.362 2 MF-it (EVA-CLIP) 0.650 0.726 0.664 0.676 0.593 0.662 259 MF-ii (EVA-CLIP) 0.649 0.723 0.665 0.670 0.592 0.660 247 InstructTA-MF-it (ours) 0.692 0.756 0.701 0.712 0.638 0.700 363 InstructTA-MF-ii (ours) 0.680 0.748 0.692 0.701 0.627 0.690 325
3.3 Dual Targeted Attack & Optimization
With the aforementioned steps, we form following dual targeted attack framework:
where $\mathcal{J}_{mf}$ is the objective of MF-it or MF-ii, following [46]. MF-it and MF-ii achieve transfer targeted attack via matching the embeddings of the adversarial example and the target response. For MF-it, its objective is defined as
[…]
where $f(\cdot)$ and $g(\cdot)$ are the paired visual encoder and text encoder of an align-based VLP model (e.g., CLIP [34], ALBEF [24], Open-CLIP [9], and EVA-CLIP [17]), respectively. For MF-ii, we have
[…]
Notably, Eqn. 6 with MF-it is marked InstructTA-MF-it, and Eqn. 6 with MF-ii is marked InstructTA-MF-ii.
When optimizing adversarial samples following Eqn. 6, this paper adopts PGD to update $\boldsymbol{x}^{\prime}$ with $T$ iterations as below,
In Algorithm 1, we outline the whole pipeline of InstructTA. Initially, once decided an adversarial target text $\boldsymbol{y}_{t}$, we use the text-to-image model $h_{\xi}$ and GPT-4 to “reverse” a target image $\boldsymbol{x}_{t}$ and produce the reference instruction $\boldsymbol{p}^{\prime}$, respectively. In each iteration, we augment the inferred instruction $\boldsymbol{p}^{\prime}$ with ChatGPT firstly. Then, we update the adversarial example with PGD over Eqn. 6. Until the end of $T$ step iterations, the final adversarial sample is obtained and used for attacking remote LVLMs.

### v3

@@ Abstract
Large vision-language models (LVLMs) have demonstrated their incredible capability in image understanding and response generation. However, this rich visual interaction also makes LVLMs vulnerable to adversarial examples. In this paper, we formulate a novel and practical targeted attack scenario that the adversary can only know the vision encoder of the victim LVLM, without the knowledge of its prompts (which are often proprietary for service providers and not publicly available) and its underlying large language model (LLM). This practical setting poses challenges to the cross-prompt and cross-model transferability of targeted adversarial attack, which aims to confuse the LVLM to output a response that is semantically similar to the attacker’s chosen target text. To this end, we propose an instruction-tuned targeted attack (dubbed InstructTA) to deliver the targeted adversarial attack on LVLMs with high transferability. Initially, we utilize a public text-to-image generative model to “reverse” the target response into a target image, and employ GPT-4 to infer a reasonable instruction $\boldsymbol{p}^{\prime}$ from the target response. We then form a local surrogate model (sharing the same vision encoder with the victim LVLM) to extract instruction-aware features of an adversarial image example and the target image, and minimize the distance between these two features to optimize the adversarial example. To further improve the transferability with instruction tuning, we augment the instruction $\boldsymbol{p}^{\prime}$ with instructions paraphrased from GPT-4. Extensive experiments demonstrate the superiority of our proposed method in targeted attack performance and transferability. The code is available at https://github.com/xunguangwang/InstructTA.
@@ 1 Introduction
Large vision-language models (LVLMs) have achieved great success in text-to-image generation [41], image captioning [9], visual question-answering [28], visual dialog [21], etc., due to the colossal increase of training data and model parameters. Benefiting from the strong comprehension of large language models (LLMs), recent LVLMs [28, 14, 57, 32, 4, 46, 2, 3, 21] on top of LLMs have demonstrated a superior capability in solving complex tasks than that of small vision-language models. It is noted that GPT-4 [37] supports visual inputs and performs human-like conversations in various complicated scenarios.
@@
However, recent studies [55, 16, 42, 5] exposed that LLMs with visual structures could be easily fooled by adversarial examples (AEs) crafted via adding imperceptible perturbations to benign images. In particular, targeted attack with malicious goals raises more serious safety concerns for LVLMs and their downstream applications. For example, targeted adversarial examples could jailbreak LVLMs to explain the bomb-making process or trick medical diagnostic models into generating reports of specific diseases for insurance. It is thus demanding to comprehensively study adversarial attacks on LVLMs to address these security threats before safety breaches happen.
@@
This paper focuses on the targeted attack on LVLMs, where the adversary aims to generate an adversarial image that misleads the LVLM to respond with a specific target text, despite varying instructions the LVLM may use. Typically, LVLMs are deployed by service providers for vertical applications in the cloud, and the built-in prompts [^1] To clarify, in this paper, we use the term “instruction” and “prompt” interchangeably. and the parameters of LLMs leveraged by the LVLMs are not publicly accessible as they are proprietary to the service providers. Conversely, it is reasonable to assume that attackers possess knowledge of the vision encoder in the LVLM, as developers often disclose the structure of the model in technical reports or attackers can acquire this information through social engineering. Consequently, adversaries can easily construct surrogate models with the same vision encoders to facilitate attacks. Nonetheless, the uncertainty surrounding the LVLM model itself and its internal instructions presents challenges for targeted attacks: (1) The attacker cannot easily gain the instruction-aware features from the vision encoder to launch exact targeted attacks due to a lack of instructions. (2) Cross-model and cross-instruction transferability places new demands on the targeted attack, since the victim LVLM may have distinct LLM backends and use varying (unknown) instructions.
@@
Figure 1: The framework of our instruction-tuned targeted attack (InstructTA). Given a target text $\boldsymbol{y}_{t}$, we first transform it into the target image $\boldsymbol{x}_{t}$ with a text-to-image model $h_{\xi}$. Simultaneously, GPT-4 infers a reasonable instruction $\boldsymbol{p}^{\prime}$. Upon providing the augmented instruction $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ which are rephrased from $\boldsymbol{p}^{\prime}$ using GPT-4, the surrogate model $M$ extracts instruction-aware features of $\boldsymbol{x}_{t}$ and the AE $\boldsymbol{x}^{\prime}$, respectively. Finally, we minimize the $L_{2}$ distance between these two features to optimize $\boldsymbol{x}^{\prime}$.
@@
  • We propose an effective targeted attack, InstructTA. It cleverly leverages LLMs to deduce reasonable prompts for target texts, leading to extracting instruction-aware features for precise targeted attacks.
  • We present the instruction-tuning paradigm in the targeted attacks to further augment the transferability of adversarial samples, benefiting from the rephrased instructions by GPT-4.
@@ 2 Related Work
LVLMs. Benefiting from the powerful capabilities of LLMs in answering diverse and complex questions [36], LVLMs [55] have demonstrated remarkable performance on a variety of multi-modal tasks, e.g., image captioning, visual question-answering, natural language visual reasoning, and visual dialog. To bridge the image-text modality gap, BLIP-2 [28] proposes the Q-Former to extract visual features into the frozen OPT [54] or FlanT5 [13]. MiniGPT-4 [57] employs the identical pre-trained ViT [18] and Q-Former as BLIP-2 but uses the more powerful Vicuna [12] to serve as the base LLM for text generation. InstructBLIP [14] introduces an instruction-aware Q-Former to extract the most task-relevant visual representations. LLaVA [32] bridges a pre-trained visual encoder with LLaMA, and fine-tunes them with question-image-answer paired data generated by GPT-4. Instead of projecting visual features onto the input embedding space of the LLM, CogVLM [46] introduces a visual expert module to synergistically incorporate image and linguistic features within the pre-trained large language model.
@@
Adversarial Attacks. Although adversarial attacks are now being developed in a variety of modalities and tasks, the earliest and most extensive research occurred in the field of image classification [6, 44]. In this case, the goal of the adversarial attack is often to fool a well-trained visual classifier by crafted adversarial images. Based on their purpose, we divide adversarial attacks into targeted attacks and non-targeted attacks [1]. Targeted attacks aim to confuse the model to recognize the adversarial example as a target label, whereas non-targeted attacks only ensure that the model makes wrong predictions, not caring about the output error category. Moreover, according to the information of the attacked model exposed to the adversary, adversarial attacks can be grouped into white-box attacks and black-box attacks. The white-box attacks can obtain the whole parameters and structure of the target model. Thus, they generally optimize adversarial perturbations with gradients of loss w.r.t inputs, e.g., FGSM [22], I-FGSM [26], C&W[7] and PGD [35]. In contrast, only the input and output of the attacked model can be captured by black-box attacks, which is more challenging and practical than white-box one. One solution is to estimate the gradient or search for successful perturbations by directly querying the target model multiple times, e.g., ZOO [10] and NES attack [25]. Another way is to implement transfer attacks by constructing adversarial samples from substitute models, e.g., SBA [39] and ensemble-based approaches [33, 17, 20, 24].
@@
AEs on LVLMs. Currently, there are some pioneering works of targeted attacks on LVLMs. On one hand, some works aim to mislead the LVLMs to output specific words. For example, Bailey et al. [5] forces the LVLM to respond with the target string in a white-box setting. Dong et al. [16] present text description attack to mislead the black-box Bard [23] to output targeted responses, which maximize the log-likelihood of the substitute LVLMs to predict the target sentence. On the other hand, some works have been devoted to fooling LVLMs into outputting text with the same semantics as the target. For example, Zhao et al. [55] propose three feature matching methods (i.e., MF-it, MF-ii and MF-tt) to launch targeted attack with given target text against LVLMs in black-box settings. Shayegani et al. [42] also align the adversarial image and target images in the embedding space for LVLM jailbreak. Different from [55], we further improve the attack transferability under the guidance of inferred and rephrased instructions. While [42] jailbreaks LVLMs with a generic prompt by perturbing the input image, our work highlights the capability to implement targeted attacks even without knowing the built-in prompts of the target LVLM.
@@ 3.1 Problem Formulation
LVLM. Given an image $\boldsymbol{x}$ and a text prompt $\boldsymbol{p}$ as inputs, the LVLM $F$ outputs an answer $\boldsymbol{y}$. Formally,
@@
where $F_{\theta}$ denotes the LVLM parameterized by $\theta$. Typically, the LVLM [28, 57, 14, 32] projects the extracted visual features into the word embedding space and then feeds them to its underlying LLM. Hence, LVLM consists of a pre-trained vision encoder, a connector and a well-trained LLM, as shown in Fig. 2.
@@
where $\boldsymbol{x}$ is a given benign image, $\boldsymbol{x}^{\prime}$ represents the adversarial example of $\boldsymbol{x}$, and $\boldsymbol{y}_{t}$ denotes the predefined target text. $\epsilon$ is the budget of controlling adversarial perturbations, and $\boldsymbol{p}$ denotes the instruction used by the target LVLM ($\boldsymbol{p}$ is unknown to attackers; see § 3.2). $S(\cdot,\cdot)$ is a metric depicting the semantic similarity between two input sentences. This way, Eqn. 2 ensures that the response of the adversarial example shall retain the same meaning as the target text $y_{t}$.
Figure 3: An example of rephrasing an instruction. Given the target response $\boldsymbol{y}_{t}$, GPT-4 infers an instruction, i.e., “Can you describe what’s in the picture you’re looking at related to food?”. The real instruction assigned to this target text is “What are the essential components depicted in this image?”
3.2 Targeted Attack with Instruction Tuning
Assumption. In this paper, we consider a practical gray-box setting, where the adversary only knows the visual encoder used in the attacked model, excluding LLMs and instructions of the victim LVLM. That is, the adversary is agnostic to $\boldsymbol{p}$. Instead, we assume that attackers know the usage domain of the target LVLM, e.g., CT image diagnosis; considering $F$ is often hosted as a task-specific service, attackers can easily obtain this information. Accordingly, attackers can locally prepare instructions $\boldsymbol{p}^{\prime}$ which are commonly used in the target domain (see below for details). We also assume that before attack, attackers can form a local surrogate model with the same vision encoder as that of the victim model.
Inferring Instructions. We first locally infer the possible instruction $\boldsymbol{p}^{\prime}$ for the target text $\boldsymbol{y}_{t}$ by using GPT-4 [37]. We design the inference prompt $\boldsymbol{p}_{g}$ as follows.
Inference prompt $\boldsymbol{p}_{g}$ What is the question or prompt for the following answer? You just provide a question or prompt. """ $\boldsymbol{y}_{t}$""
According to the above prompt, we can derive a reasonable instruction $\boldsymbol{p}^{\prime}$ by feeding the target response text $\boldsymbol{y}_{t}$ to GPT-4.
Forming Surrogate Models. The surrogate model comprises two modules: the visual encoder sourced from the victim model and the pre-trained projector. The projector follows the visual encoder, as depicted in Fig. 1. By providing an instruction as input, the projector is capable of mapping the existing global features into informative features that are specifically tailored to the given instruction. In this paper, Q-Former [28] is selected as the projector. Consequently, once the instructions have been inferred, the surrogate model employs the projector to convert the features obtained from the visual encoder into instruction-aware features.
Augmenting Transferability. Our tentative exploration shows that AEs generated by Eqn. 3, although manifesting reasonable attackability, are often less transferable to the wide range of possible instructions in the wild. Thus, to improve the transferability and robustness of the generated adversarial examples, we use GPT-4 to rephrase the inferred $\boldsymbol{p}^{\prime}$ as augmentation. The prompt $\boldsymbol{p}_{r}$ for paraphrasing is formulated as follows.
Rephrasing prompt $\boldsymbol{p}_{r}$ Rephrase the following sentence without changing its original meaning. Please give $n-1$ examples of paraphrasing that look different and begin the statement with "-". """ $\boldsymbol{p}^{\prime}$""
where $n-1$ represents the number of sentences rephrased by GPT-4. Together with the original inferred instruction $\boldsymbol{p}^{\prime}$, a total of $n$ instructions are collected in a set $Q$. We set $n$ to 10 by default. Benefiting from the augmented instructions, Eqn. 3 can be rewritten as
$\min_{\boldsymbol{x}^{\prime}}\mathcal{J}_{ins}=\mathbb{E}_{\boldsymbol{p}_{i}^{\prime},\boldsymbol{p}_{j}^{\prime}\in Q}\|M(\boldsymbol{x}^{\prime}$ $,\boldsymbol{p}_{i}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}_{j}^{\prime})\|_{2}^{2},$ (4)
[…]
where $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ are randomly selected from the set $Q$. We utilize the rephrasing prompt $\boldsymbol{p}_{r}$ to construct the instruction set $Q$ via GPT-4. As illustrated in Fig. 3, GPT-4 is generally effective to rephrase a seed instruction $\boldsymbol{p}^{\prime}$ into a set of semantically-correlated instructions.
3.3 Dual Targeted Attack & Optimization
With the aforementioned steps, we form the following dual targeted attack framework:
where $\mathcal{J}_{mf}$ is the objective of MF-it, following [55]. MF-it achieves transfer targeted attack via matching the embeddings of the adversarial example and the target response. For MF-it, its objective is defined as
[…]
where $f(\cdot)$ and $g(\cdot)$ are the paired visual encoder and text encoder of an align-based VLP model (e.g., CLIP [40], ALBEF [29], Open-CLIP [11], and EVA-CLIP [19]), respectively.
When optimizing adversarial samples following Eqn. 5, this paper adopts PGD to update $\boldsymbol{x}^{\prime}$ with $T$ iterations as below,
In Algorithm 1, we outline the whole pipeline of InstructTA. Initially, once decided an adversarial target text $\boldsymbol{y}_{t}$, we use the text-to-image model $h_{\xi}$ and GPT-4 to “reverse” a target image $\boldsymbol{x}_{t}$ and produce the reference instruction $\boldsymbol{p}^{\prime}$, respectively. Furthermore, we augment the inferred instruction $\boldsymbol{p}^{\prime}$ with GPT-4 to form the instruction set $Q$. In each iteration, we randomly select instructions $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ from $Q$. Then, we update the adversarial example with PGD over Eqn. 5. Until the end of $T$ step iterations, the final adversarial sample is obtained and used for attacking remote LVLMs.
@@ 3.3 Dual Targeted Attack & Optimization
0:  a benign image $\boldsymbol{x}$, a target text $\boldsymbol{y}_{t}$, a text-to-image generative model $h_{\xi}$, a surrogate model $M$, a VLP model with the visual encoder $f$ and text encoder $g$, GPT-4 $G_{4}$, custom prompts $\boldsymbol{p}_{g}$ and $\boldsymbol{p}_{r}$, step size $\eta$, perturbation budget $\epsilon$, attack iterations $T$
1:  Initialize: $\boldsymbol{x}^{\prime}=\boldsymbol{x}$, $\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$,   // generate a target image $\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t})$,  //infer an instruction $Q\leftarrow G_{4}(\boldsymbol{p}_{r}\boldsymbol{p^{\prime}})$  // build an instruction set $Q$ by rephrasing $\boldsymbol{p^{\prime}}$
@@
3:     $\boldsymbol{p}_{i}^{\prime}\sim Q$, $\boldsymbol{p}_{j}^{\prime}\sim Q$  // randomly select instructions
4:     $\mathcal{L}=\mathcal{J}_{ins}-\mathcal{J}_{mf}$  // Eqn. 5